Q1In the circuit given below, R = 52. Find Vy and in 0.8i. 208 v 5 A 102 0.25v, 3A The value of vy in the circuit is V. The val

Q2Determine the gain vol vs of the given transistor amplifier circuit where / = 45,- 210 4 kl 1 WW * 4 ke 1000+ 10k22 The value
please i want the right answer for them
Answer #1

- R=52, Vx = ? lo=? + 203 vx + SA {lor vi 0. v 50g - 1 022 Os T . Th T 3+0.250x = - ve - 3 = -O-SV1-0.25V1 - 3 -0.75VX ve = -4K M Mlo 2K b12 A 468 A 1000 I vo slok Vs I=4500 (Given) Vo = -I (DK) = - 4510xlok IVO = - 4Go X103 100 Iz= 10 . mesh analysiBy equation & - Vs = 6x103 (2.1125 Iz) - 4x103 I 2 = 12.675 X103 I₂ - 4 x 10 Iz us = 8.675 x 10²127 D By equation ♡ & Vo = -in this we used basic kvl kcl to find voltage, current and Voltage gain.